The formula A=127e0.034t models the population of Corona, CA, in thousands, t years after 1994. When will the population of the city reach 175 thousand?
Answer #1

A = 0-0340 127 e when A = 175 thousand .: A = 175 0.034t 127 e 175 +0.0345 e 175 127 Leena x = lna @.0341 en 199) ta las 173/